### v1

@@ Abstract
With the rapid development of machine learning in recent years, many problems in meteorology can now be addressed using AI models. In particular, data-driven algorithms have significantly improved accuracy compared to traditional methods. Meteorological data is often transformed into 2D images or 3D videos, which are then fed into AI models for learning. Additionally, these models often incorporate physical signals, such as temperature, pressure, and wind speed, to further enhance accuracy and interpretability. In this paper, we review several representative AI + Weather/Climate algorithms and propose a new paradigm where observational data from different perspectives, each with distinct physical meanings, are treated as multimodal data and integrated via transformers. Furthermore, key weather and climate knowledge can be incorporated through regularization techniques to further strengthen the model’s capabilities. This new paradigm is versatile and can address a variety of tasks, offering strong generalizability. We also discuss future directions for improving model accuracy and interpretability.
keywords:
Weather and Climate Sciences, AI, Hybrid Models, Multi-modality, Transformers
@@ 1 Introduction
The field of meteorological science, encompassing both weather and climate sciences, presents some of the most intricate and dynamically evolving challenges within contemporary scientific research [1, 2]. Accurately understanding and predicting the behavior of atmospheric systems, ocean currents, and climate variables demands the analysis of massive datasets and the integration of diverse observational models. These challenges stem from the inherent variability of weather and climate phenomena, their long-term temporal scales, and the spatially heterogeneous nature of the data involved [3, 4, 5]. In light of these complexities, the role of AI technologies in meteorological science has become increasingly critical. Traditional numerical models, often grounded in deterministic physical equations, face significant limitations, underscoring the potential of data-driven AI methods to enhance and expand the accuracy and scope of scientific investigations in this field.
@@
Machine learning, and in particular deep learning [6], has brought significant advantages to meteorological science [7]. The ability to leverage large-scale computational resources, such as GPUs, has led to breakthroughs in processing and analyzing complex datasets efficiently and effectively. Deep learning-based models [8, 9, 10] can automate many processes that previously require much manual effort. These methods open up possibilities for more accurate simulations and predictions in weather and climate modeling, where even small improvements in accuracy can lead to profound real-world impact, from policy decisions to disaster response.
However, despite these promising advances, there are inherent difficulties when applying purely data-driven approaches to weather and climate science. The vastness of the input data—spanning multiple dimensions such as time, space, and atmospheric variables—introduces the risk of overfitting or inadequate model convergence. A model that learns solely from observed data may lack the physical constraints that govern atmospheric processes. Without these constraints, models may produce results that are mathematically plausible but physically unrealistic. To address these limitations, researchers have increasingly turned to hybrid models that combine data-driven approaches with established physical principles [8, 11, 12]. By embedding physical constraints into the training process or combining neural networks with traditional numerical models, hybrid approaches ensure that the outputs are both scientifically valid and computationally efficient.
Although showing great potential for advancing weather and climate sciences, current hybrid models still face several critical limitations. The lack of a unified architecture complicates the design and optimization process, hindering standardization and comparability across studies. Limited multi-modal data fusion reduces their ability to fully capture the complexity of natural systems. Moreover, challenges in scaling these models to large, complex systems further restrict their practical application. Addressing these challenges will require more standardized frameworks, enhanced data integration techniques, and improved methods for handling complex systems, ultimately enabling hybrid models to achieve better accuracy, scalability, and reliability.
This paper introduces a standardized framework for integrating heterogeneous observational data, each with distinct physical interpretations, into a cohesive model. We propose a physics-guided multimodal approach, leveraging transformer architectures, specifically tailored for weather and climate sciences. This framework serves as a versatile foundation for addressing a broad spectrum of meteorology-related challenges, ranging from extreme weather prediction to long-term climate projections. Transformers have demonstrated considerable efficacy in processing sequential and structured data, making them particularly well-suited for time-series forecasting and modeling spatiotemporal dynamics in climate systems. In a multimodal context, transformers can seamlessly integrate a variety of data sources—ranging from satellite imagery and weather station measurements to atmospheric probe sensor data—into a unified model. Additionally, we examine the incorporation of critical weather and climate knowledge in three distinct ways, showing how this integration can significantly enhance both predictive accuracy and model interpretability. This approach not only improves the precision of climate predictions but also contributes to a deeper understanding of the fundamental physical mechanisms driving weather and climate phenomena.
@@
2 Categories of AI Models in Weather and Climate Sciences
2.1 Forecasting Models
Forecasting models in weather and climate sciences predict future meteorological conditions using historical data. Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs) have been widely used for their ability to capture spatial dependencies and model complex meteorological relationships. CNNs are effective in identifying spatial patterns in variables like temperature and pressure, while GNNs handle dynamic, grid-based weather system data across regions [13, 14]. For instance, MetNet combines CNNs, LSTMs, and autoencoders to capture multi-scale spatial and temporal features for short-term forecasts [13]. Similarly, GraphCast employs GNNs within an ”encode-process-decode” framework to model intricate spatiotemporal relationships, enhancing flexibility and accuracy over traditional approaches [14].
However, CNN and GNN-based models often fail to capture long-range temporal dependencies and global interactions crucial for accurate climate forecasting. Transformer-based models, leveraging attention mechanisms, address these limitations by focusing on relevant features across space and time [15, 16, 17]. Notable examples include PanGu Weather [15], which uses a 3D-Swin Transformer for spatiotemporal data processing, and FourCastNet [16], integrating adaptive Fourier neural operators with vision Transformers for high-resolution data. These models better capture complex dependencies and model both local and global features effectively. These developments indicate a shift towards more scalable and flexible solutions, improving predictions of extreme weather events and long-term climate changes.
To optimize these models, we examine the classical objective function used in forecasting tasks. For example, the objective for PanGu Weather [15] is:
$\mathcal{L}=\text{MSE}(X_{t+1},f(X_{t},\theta)),$ (1)
$X_{t+1}\sim P(X_{t+1}|X_{1},\cdots,X_{t}),$ (2)
2.2 Classification and Detection Models
Classification and detection models in weather and climate sciences identify specific phenomena, such as extreme weather events, by classifying or localizing occurrences like storms and heatwaves. These models improve localization accuracy and address limited labeled data by extracting meaningful patterns from historical data for future predictions [18, 19, 20]. Deep learning architectures, particularly CNNs, have been widely adopted for climate pattern detection and extreme event classification [18, 19]. Several works combining CNNs with recurrent structures like LSTMs capture the spatio-temporal dynamics of events such as typhoons [20].
$\mathcal{L}=-\sum_{i=1}^{N}\overline{y}_{t}\log(f(X_{t},\theta)),$ (3)
These tasks can be generalized as next-token prediction problems:
$Y_{t}\sim P(Y_{t}\mid X_{1},\cdots,X_{t}),$ (4)
2.3 Re-scaling Models
Re-scaling techniques, especially super-resolution (SR), enhance the resolution and quality of meteorological data by generating high-resolution outputs from lower-resolution inputs, crucial for accurate climate modeling [21, 22, 23, 24]. Deep learning models, particularly CNNs, effectively learn mappings between low and high-resolution data, capturing spatial and temporal patterns to improve dataset granularity. A primary application is downscaling precipitation data from coarse (e.g., 100 km) to fine resolutions (e.g., 12.5 km), better representing localized phenomena like extreme events and supporting more accurate weather predictions and climate assessments.
The objective function for re-scaling tasks typically uses MSE:
$\mathcal{L}=\text{MSE}(X_{\text{HR}},f(X_{\text{LR},\theta})),$ (5)
where $X_{\text{HR}_{t}}$ is the high-resolution data at time $t$.
More generally, re-scaling can be expressed as:
$Y_{t}\sim P(Y_{t}\mid X_{1},\cdots,X_{t}),$ (7)
2.4 Universal Embedding
Universal embedding techniques represent diverse meteorological data modalities within a common space, enabling integration across various meteorological-related tasks. This unified approach enhances model flexibility and adaptability [25, 26]. Deep learning architectures, such as LSTMs and Masked AutoEncoders, have proven effective in creating universal embeddings. For example, Kratzert et al. [25] used EA-LSTM to model rainfall–runoff processes, while W-MAE [26] employed self-supervised pre-training on ERA5 datasets, improving meteorological variable predictions and enhancing models like FourCastNet.
The objective for universal embedding is:
where $E_{t}$ is the embedding at time $t$, and $[X,Y,Z,\cdots]_{1},\cdots,[X,Y,Z,\cdots]_{t}$ are inputs from various modalities. This framework integrates multiple data sources into a unified space, supporting diverse tasks like forecasting, classification, and anomaly detection.
2.5 Relationship Mining
Relationship mining in climate analysis explores connections between climatic factors, essential for understanding climate dynamics. Early methods used graph-based techniques to identify teleconnections [27], while recent studies employ GNNs for more effective modeling [28, 29, 30].
$G^{\prime}_{t}\sim P(G^{\prime}_{t}\mid G_{1},G_{2},\cdots,G_{t})$ (10)
2.6 Others
3 State-of-the-art Hybrid Models
@@
While AI models, especially those based on deep learning, have shown promise in weather forecast and climate prediction, their reliance on data alone presents significant challenges. These models can produce accurate results, but often suffer from issues related to poor interpretability, limited generalization, and an inability to fully grasp the underlying physical laws governing the atmosphere. In contrast, hybrid models [8], which combine both deep learning and physical modelling, offer a powerful solution to these challenges. By integrating the strengths of data-driven approaches with the robustness of theory-based models, hybrid approaches can leverage the flexibility and adaptability of machine learning while maintaining the interpretability and extrapolation capabilities of physical models.
@@
Regularization through Loss or Output CM2Mc-LPJmL ESM [11] Earth system models (ESM) simulation Enforcing a physical constraint on outputs to preserve global precipitation sums
NeuralGCM [31] Physical simulation of Earth’s atmosphere Using a combination of three loss functions to encourage accuracy and sharpness while penalizing bias.
DeepHybrid [32] Weather Forecasting Applying physical laws constraints on the output to be spatially smooth.
Utilizing Existing Physical Models (PDEs) NowcastNet [12] Extreme precipitation nowcasting Combining physical-evolution schemes and conditional-learning methods
ClimODE [33] Climate and weather prediction Learning a continuous-time PDE model, grounded on continuity equation
Physics as an Input Signal CLLMate [34] Weather and Climate Events Forecasting Knowledge graph and meteorological raster data as inputs of multimodal LLM
DUQ [35], ECMWF Enhancement [36] Weather Forecasting Incorporating prior knowledge from NWP with information fusion mechanism.
@@ 3.1 Regularization through Loss or Output
General Circulation Models (GCMs) serves as the foundation for simulating and predicting atmospheric dynamics. NeuralGCM [31] is a hybrid model for weather and climate by combining GCMs and machine learning. Its inputs include the prognostic variables $X_{1},\cdots,X_{t}$, and other necessary variables $V$ like forcings and noise. NueralGCM will output the prognostic variable tendencies $X_{t+1}$ in the future. The process falls in the scope of Eq. 2: $X_{t+1}\sim P(X_{t+1}|X_{1},\cdots,X_{t},V)$. The training object for NueralGCM is a combination of three loss functions. In addition to the standard MSE loss on the main predicted variables, a regularization to force the spectrum to match the training data, and a MSE loss term on the spherical harmonic coefficient’s mean magnitude. In general, if the AI model predicts unrealistic velocity profiles or deviates from established physical behaviors like conservation of mass or momentum, a regularization term can be used to adjust the predictions back to plausible values, improving both the accuracy and physical fidelity of the model.
@@
In climate modeling, particularly for precipitation, models aim to capture the complex dynamics of moisture transfer, cloud formation, and precipitation processes. If a model’s predictions for precipitation are unrealistic or violate known physical principles (e.g., precipitation gradients that are overly smooth or unrealistic in regions where intense variability is expected), a physical constraint can be incorporated as an additional layer in the generator network architecture after training to ensure the preservation of the global precipitation sum [11]. The post-processing is performed by a generator, which is a simplification form of Eq. 7: $X^{\prime}_{t}\sim P(X^{\prime}_{t}|X_{t})$. A physical constraint to preserve the total global precipitation amount of CM2Mc-LPJmL model is applied on the output:
@@ 3.2 Utilizing Existing Physical Models
The physical principles that are widely used in climate models include the law of conservation of mass, the ideal gas law, the first law of thermodynamics and continuity equations [7]. A powerful method for integrating physics into AI models is to incorporate existing physical models such as partial differential equations (PDEs) that describe fundamental physical processes into the model’s architecture. This approach can help the AI model to better respect the underlying dynamics of the physical system it is modeling.
@@
where $u$ is a quantity such as temperature driven by the flow’s velocity $v$ and sources $s$. AI models can be integrated with these equations by using Physics-Informed Neural Networks (PINNs), where the neural network learns to approximate solutions to the PDEs. For example, ClimODE [33] introduces a neural network to predict flow velocity and utilize the predicted velocity in the continuity equation for weather prediction. It is a typical application of Eq. 2: $X_{t+1}\sim P(X_{t+1}|X_{1},\cdots,X_{t})$, where $X$ is the quantity variables to be predicted, and $P()$ consists of flow velocity network and the continuity equation (PDE in Eq. 12) is solved to obtain the prediction results. By incorporating PDE equation as part of the model’s architecture or training process, the AI model can ensure its predictions are physically consistent with how quantities should propagate through the environment. This approach allows the AI to use the known mathematical description of the process to guide its learning, resulting in more accurate and realistic predictions of quantities changes over time.
@@ 3.3 Physical Knowledge as an Input Signal
In weather forecast, key physical priors like weather and climate knowledge, as well as common-sense knowledge are used as input features for AI models. For instance, in CLLMate [34], a multimodal large language model (LLM) for weather and climate event prediction, the authors propose a framework that uses a LLM to extract information from environment-focused news articles and build a knowledge graph. This graph, combined with meteorological raster data, is used in the multimodal LLM to predict weather and climate events. By incorporating such input signals into the model, it can learn not just from the raw observational data (e.g., satellite images, sensor data), but also from the underlying physical processes that govern weather dynamics. It can be viewed as an example of Eq. 4: $Y_{t}\sim P(Y_{t}|X_{1},\cdots,X_{t},S)$, where $X$ is the raw observational data, $S$ is the physical knowledge, and $Y_{t}$ is weather and climate event prediction results. The multimodal LLM with parameters $\Theta$ performs the whole prediction process.
@@ 4.3 Challenges in Scaling to Large, Complex Systems
5 Perspective: Towards a Unified Multimodal Transformer Path for AI4Weather/Climate
In this work, we propose a novel approach to advancing climate science through the integration of a physics-guided multimodal transformer framework. This approach aims to address the challenges associated with existing hybrid models by seamlessly combining diverse weather and climate data modalities while embedding critical physical principles. By doing so, we highlight how transformer models can not only handle complex, multi-dimensional atmospheric data but also leverage physical knowledge to enhance model interpretability, accuracy, and scientific validity.
@@ 5.1 Multi-modal Fusion
  • For image-like data $X_{I}\in\mathbb{R}^{H\times W}$, a transformer encoder (e.g., ViT [37]) is employed to extract semantic representations, which are then flattened into a sequence of tokens $X^{\prime}_{I}\in\mathbb{R}^{N\times D}$.
  • For video-like data $X_{V}\in\mathbb{R}^{T\times H\times W}$, the video can be converted into spacetime patches by compressing both the spatial and temporal dimensions into a latent space [38]. These patches are subsequently flattened into a sequence of tokens $X^{\prime}_{V}\in\mathbb{R}^{N\times D}$ for further processing.
  • When dealing with time series data $X_{S}\in\mathbb{R}^{T}$, the data can be divided into overlapping or non-overlapping patches [39], which are then linearly projected into the target domain $X^{\prime}_{S}\in\mathbb{R}^{N\times D}$.
  • For text data $X_{T}\in\mathbb{N}^{T}$, a tokenizer such as BBPE [40] is used to convert the text into discrete tokens $X^{\prime}_{T}\in\mathbb{R}^{N\times D}$.

### v2

@@ Abstract
This position paper argues that the next generation of artificial intelligence in meteorological and climate sciences must transition from fragmented hybrid heuristics toward a unified paradigm of physics-guided multimodal transformers. While purely data-driven models have achieved significant gains in predictive accuracy, they often treat atmospheric processes as mere visual patterns, frequently producing results that lack scientific consistency or violate fundamental physical laws. We contend that current “hybrid” attempts to bridge this gap remain ad-hoc and struggle to scale across the heterogeneous nature of meteorological data ranging from satellite imagery to sparse sensor measurements. We argue that the transformer architecture, through its inherent capacity for cross-modal alignment, provides the only viable foundation for a systematic integration of domain knowledge via physical constraint embedding and physics-informed loss functions. By advocating for this unified architectural shift, we aim to steer the community away from “black-box” curve fitting and toward AI systems that are inherently falsifiable, scientifically grounded, and robust enough to address the existential challenges of extreme weather and climate change.
Machine Learning, ICML
@@ 1 Introduction
The field of meteorological science, encompassing both weather and climate sciences, presents some of the most intricate and dynamically evolving challenges within contemporary scientific research (Edwards, 2011; Stute et al., 2001). Accurately understanding and predicting the behavior of atmospheric systems, ocean currents, and climate variables demands the analysis of massive datasets and the integration of diverse observational models. These challenges stem from the inherent variability of weather and climate phenomena, their long-term temporal scales, and the spatially heterogeneous nature of the data involved (Bauer et al., 2015; Abraham et al., 2002).
@@
In light of these complexities, machine learning, and in particular deep learning (LeCun et al., 2015), has brought transformative advantages to meteorological science (Ren et al., 2021). The ability to leverage large-scale computational resources has led to breakthroughs in processing complex datasets, opening possibilities for more accurate simulations where even incremental improvements can lead to profound real-world impacts, from policy decisions to disaster response.
However, we argue that the current trajectory of AI-driven meteorology is reaching a critical crossroads. Purely data-driven approaches, while high in predictive accuracy, often treat atmospheric variables as mere numerical pixels, ignoring the fundamental physical constraints that govern our planet. To address this, the community has turned to hybrid models that combine neural networks with established physical principles (Reichstein et al., 2019; Hess et al., 2022). Yet, we observe that many current hybrid approaches remain fragmented and ad-hoc, typically relying on task-specific heuristics or simple feature-concatenation. This lack of a unified architecture hinders the field’s ability to systematically scale across heterogeneous data sources, such as satellite imagery, sensor networks, and numerical outputs, into a coherent, scientifically consistent model.
This position paper argues that the future of meteorological AI must move beyond fragmented hybrid heuristics toward a unified paradigm: Physics-Guided Multimodal Transformers. We contend that the transformer architecture is not merely another neural backbone, but a unique mathematical bridge capable of aligning heterogeneous observational data with explicit domain knowledge. We advocate for a shift toward frameworks that systematically embed physical laws into the multi-head attention mechanisms and architectural design, rather than treating physical signals as secondary inputs.
By adopting this multimodal transformer paradigm, the field can transcend “black-box” curve fitting and develop foundational tools for next-generation meteorological systems. Such systems will be capable of addressing diverse challenges ranging from extreme weather forecasting to climate change projection with improved reliability, falsifiability, and scientific consistency. In the following sections, we outline three distinct integration mechanisms and identify the necessary steps for the community to realize this unified vision.
@@
2 State-of-the-art Hybrid Models in Weather and Climate Sciences
@@
While AI models, especially those based on deep learning, have shown promise in weather forecast and climate prediction, their reliance on data alone presents significant challenges. These models can produce accurate results, but often suffer from issues related to poor interpretability, limited generalization, and an inability to fully grasp the underlying physical laws governing the atmosphere. In contrast, hybrid models (Reichstein et al., 2019), which combine both deep learning and physical modelling, offer a powerful solution to these challenges. By integrating the strengths of data-driven approaches with the robustness of theory-based models, hybrid approaches can leverage the flexibility and adaptability of machine learning while maintaining the interpretability and extrapolation capabilities of physical models.
@@
Regularization through Loss or Output CM2Mc-LPJmL ESM (Hess et al., 2022) Earth system models (ESM) simulation Enforcing a physical constraint on outputs to preserve global precipitation sums
NeuralGCM (Kochkov et al., 2024) Physical simulation of Earth’s atmosphere Using a combination of three loss functions to encourage accuracy and sharpness while penalizing bias.
DeepHybrid (Grover et al., 2015) Weather Forecasting Applying physical laws constraints on the output to be spatially smooth.
Utilizing Existing Physical Models (PDEs) NowcastNet (Zhang et al., 2023) Extreme precipitation nowcasting Combining physical-evolution schemes and conditional-learning methods
ClimODE (Verma et al., 2024) Climate and weather prediction Learning a continuous-time PDE model, grounded on continuity equation
Physics as an Input Signal CLLMate (Li et al., 2024) Weather and Climate Events Forecasting Knowledge graph and meteorological raster data as inputs of multimodal LLM
DUQ (Wang et al., 2019), ECMWF Enhancement (Frnda et al., 2019) Weather Forecasting Incorporating prior knowledge from NWP with information fusion mechanism.
@@ 2.1 Regularization through Loss or Output
General Circulation Models (GCMs) serves as the foundation for simulating and predicting atmospheric dynamics. NeuralGCM (Kochkov et al., 2024) is a hybrid model for weather and climate by combining GCMs and machine learning. Its inputs include the prognostic variables $X_{1},\cdots,X_{t}$, and other necessary variables $V$ like forcings and noise. NueralGCM will output the prognostic variable tendencies $X_{t+1}$ in the future. The process falls in the scope of Eq. 7: $X_{t+1}\sim P(X_{t+1}|X_{1},\cdots,X_{t},V)$. The training object for NueralGCM is a combination of three loss functions. In addition to the standard MSE loss on the main predicted variables, a regularization to force the spectrum to match the training data, and a MSE loss term on the spherical harmonic coefficient’s mean magnitude. In general, if the AI model predicts unrealistic velocity profiles or deviates from established physical behaviors like conservation of mass or momentum, a regularization term can be used to adjust the predictions back to plausible values, improving both the accuracy and physical fidelity of the model.
@@
In climate modeling, particularly for precipitation, models aim to capture the complex dynamics of moisture transfer, cloud formation, and precipitation processes. If a model’s predictions for precipitation are unrealistic or violate known physical principles (e.g., precipitation gradients that are overly smooth or unrealistic in regions where intense variability is expected), a physical constraint can be incorporated as an additional layer in the generator network architecture after training to ensure the preservation of the global precipitation sum (Hess et al., 2022). The post-processing is performed by a generator, which is a simplification form of Eq. 12: $X^{\prime}_{t}\sim P(X^{\prime}_{t}|X_{t})$. A physical constraint to preserve the total global precipitation amount of CM2Mc-LPJmL model is applied on the output:
@@ 2.2 Utilizing Existing Physical Models
The physical principles that are widely used in climate models include the law of conservation of mass, the ideal gas law, the first law of thermodynamics and continuity equations (Ren et al., 2021). A powerful method for integrating physics into AI models is to incorporate existing physical models such as partial differential equations (PDEs) that describe fundamental physical processes into the model’s architecture. This approach can help the AI model to better respect the underlying dynamics of the physical system it is modeling.
@@
where $u$ is a quantity such as temperature driven by the flow’s velocity $v$ and sources $s$. AI models can be integrated with these equations by using Physics-Informed Neural Networks (PINNs), where the neural network learns to approximate solutions to the PDEs. For example, ClimODE (Verma et al., 2024) introduces a neural network to predict flow velocity and utilize the predicted velocity in the continuity equation for weather prediction. It is a typical application of Eq. 7: $X_{t+1}\sim P(X_{t+1}|X_{1},\cdots,X_{t})$, where $X$ is the quantity variables to be predicted, and $P()$ consists of flow velocity network and the continuity equation (PDE in Eq. 2) is solved to obtain the prediction results. By incorporating PDE equation as part of the model’s architecture or training process, the AI model can ensure its predictions are physically consistent with how quantities should propagate through the environment. This approach allows the AI to use the known mathematical description of the process to guide its learning, resulting in more accurate and realistic predictions of quantities changes over time.
@@ 2.3 Physical Knowledge as an Input Signal
In weather forecast, key physical priors like weather and climate knowledge, as well as common-sense knowledge are used as input features for AI models. For instance, in CLLMate (Li et al., 2024), a multimodal large language model (LLM) for weather and climate event prediction, the authors propose a framework that uses a LLM to extract information from environment-focused news articles and build a knowledge graph. This graph, combined with meteorological raster data, is used in the multimodal LLM to predict weather and climate events. By incorporating such input signals into the model, it can learn not just from the raw observational data (e.g., satellite images, sensor data), but also from the underlying physical processes that govern weather dynamics. It can be viewed as an example of Eq. 9: $Y_{t}\sim P(Y_{t}|X_{1},\cdots,X_{t},S)$, where $X$ is the raw observational data, $S$ is the physical knowledge, and $Y_{t}$ is weather and climate event prediction results. The multimodal LLM with parameters $\Theta$ performs the whole prediction process.
@@ 3.3 Challenges in Scaling to Large, Complex Systems
4 Perspective: Towards a Unified Multimodal Transformer Path
In this work, we propose to advance climate science through the integration of a physics-guided multimodal transformer framework. This approach aims to address the challenges associated with existing hybrid models by seamlessly combining diverse weather and climate data modalities while embedding critical physical principles. By doing so, we highlight how transformer models can not only handle complex, multi-dimensional atmospheric data but also leverage physical knowledge to enhance model interpretability, accuracy, and scientific validity.
@@ 4.1 Multi-modal Fusion
  • For image-like data $X_{I}\in\mathbb{R}^{H\times W}$, a transformer encoder (e.g., ViT (Dosovitskiy et al., 2021)) is employed to extract semantic representations, which are then flattened into a sequence of tokens $X^{\prime}_{I}\in\mathbb{R}^{N\times D}$.
  • For video-like data $X_{V}\in\mathbb{R}^{T\times H\times W}$, the video can be converted into spacetime patches by compressing both the spatial and temporal dimensions into a latent space (Brooks et al., 2024). These patches are subsequently flattened into a sequence of tokens $X^{\prime}_{V}\in\mathbb{R}^{N\times D}$ for further processing.
  • When dealing with time series data $X_{S}\in\mathbb{R}^{T}$, the data can be divided into overlapping or non-overlapping patches (Nie et al., 2023), which are then linearly projected into the target domain $X^{\prime}_{S}\in\mathbb{R}^{N\times D}$.
  • For text data $X_{T}\in\mathbb{N}^{T}$, a tokenizer such as BBPE (Wang et al., 2020) is used to convert the text into discrete tokens $X^{\prime}_{T}\in\mathbb{R}^{N\times D}$.